OpenStudy (anonymous):

Terrible exponents Compute A divided by B where A = 66666^4 and B = 22222^4 How do I start? @e.mccormick

OpenStudy (e.mccormick):

If you think about the multiplication definition of an exponent and factoring, it should be eaier to do.

OpenStudy (anonymous):

Umm, what do you mean by that?

OpenStudy (e.mccormick):

\(a^2=a\cdot a\) right? and \(\dfrac{a^2}{a}=\dfrac{a\cdot a}{a}=\dfrac{\cancel{a}\cdot a}{\cancel{a}}=a\)

OpenStudy (anonymous):

Hmm...

OpenStudy (e.mccormick):

So what could you factor out of your problem and cancel?

OpenStudy (anonymous):

22222?

OpenStudy (e.mccormick):

Well, 22222 or \(22222 ^4\) Both are possible.

OpenStudy (anonymous):

 How do I do that?